# Learning from Emergence: A Study on Proactively Inhibiting the Monosemantic Neurons of Artificial Neural Networks

### v2

@@ 4.1. Experimental Setup
Table 2. The experimental results of Swin-Transformer on the ImageNet data and ConvGRU on the data HKO-7. For results on ImageNet-1k dataset, 3 Swin-Transformers pretrained on ImageNet-22k are used as backbones. The metric used is top-1 accuracy, where a higher value indicates better performance. For results on HKO-7 dataset, we initially train a ConvGRU model for 20k steps to create the base model. The metrics used are B-MSE and B-MAE, where a smaller value indicates better performance. The best results are in bold fonts.
Model Swin-T Swin-S Swin-B B-MAE B-MSE
@@
Original 80.9 83.2 85.1 1003.41 309.96
Naive (a) 81.0 83.4 84.6 1003.56 309.83
Naive (b) 81.0 83.4 85.1 1003.40 310.13
MEmeL 81.1 83.4 85.1 1003.25 209.94
MEmeL-Tune 81.1 83.5 85.2 998.81 298.16
@@ 4.1.1. Data Sets, Base Models, and Tasks
The configuration of above base models generally follow their original settings (see more details in Appx. E.1). The metrics can be found in Appx. E.2. We emphasize that our method is a general module applicable to models of any scale. After applying our module, the effectiveness of MEmeL can be evaluated by comparing it with the original model.
@@ 4.1.2. Hyper-parameter Setting
This paper focuses on the performance improvement that is brought by introducing reverse deactivation into milestone models. To fairly validate the effectiveness of the proposed MEmel, we deactivate the neuron with only top-$1$ MS in each batch (recorded as $\rm{MEmeL}$). In other words, we give up the way of determining the number of neurons to be deactivated through hyper-parameter tuning. Therefore, setting may somehow compromise the score improvement, but it preserves fairness. Besides, we also display the results with tuned hyper-parameters (recorded as $\rm{MEmeL-Tune}$). See more details in Appx. E.3.
@@ 4.1.3. Implementation
All the codes are implemented in PyTorch (Adam et al., 2017), which is available in the supplementary materials. Our experiment is conducted on 4 V100 GPU cards. Our code are available through the anonymous repo.
@@ 4.2. Main Experiment Result
Tab. 1 and Tab. 2 demonstrate the performance of MEmeL incorporated with BERT, Swin-Transformer, and ConvGRU, respectively. The “Original” method denotes the raw model, “Naive (a)” and “Naive (b)” are corresponding to the original model incorporated with naive inhibition methods (see details in Eq. (10)). By comparing the original method and the method enhanced with MEmeL, we can see that the MEmeL (especially MEmeL-Tune) achieves better or comparable results on different tasks, different data sets, and different base models. This validates that neural networks can achieve better results by deactivating monosemantic neurons. By comparing Naive (a), (b) with MEmeL (and MEmeL-Tune), we can easily observe that naive deactivation methods may intensify the monosemanticity of neural networks, which may lead to inferior performance improvement as discussed in Sec. 3.2.1.
@@ 4.3. Case Study on Deactivating Monosemantic Neurons
To validate that the monosemanticity is indeed inhibited by our reverse deactivation, we conducted experiments on the ImageNet-1k dataset using the Swin-Transformer model as shown in Tab. 3. We collected 10k samples for 4 different settings, where the modification of $h$ was done using Original (no modification), and two naive methods (a) and (b), and Reverse Deactivation in Sec. 3.2.2.
@@ 4.4. Potential and Limitation of MEmeL
According to our hypothesis, MEmeL induces the model to accumulate general and abstract functionality instead of monosemanticity for a specific task, which is consistent with the goal of per-taining. Although MEmeL achieves good results during fine-tuning (demonstrated at Main Experiments in Sec. 4.2), the improvement is expected to be even greater when it is applied to the pre-training phase.
The main obstacle is the high computational resource cost required for per-taining. Currently, we have only completed validation on relatively small precipitation forecasting datasets. In Tab. 4, the model pre-trained with MEmeL (P-MEmeL) outperforms the one without it (P-Original). Based on these two models, the same finetuning process is conducted without MEmeL. Using MEmeL during pretraining improves B-MAE by 0.18% and B-MSE by 1.29% (T-MEmeL). In contrast, using MEmeL during finetuning improves B-MAE by 0.02% B-MSE by 0.01% (MEmeL in Table. 2). The results validate our hypothesis.
@@ Appendix A An Example Analogy that Helps Illustrate Our Motivation
From the perspective of scale differences, when the model is small, the limited number of neurons makes it difficult to abstract and analyze the input, and to provide functional support for integrating a large amount of scattered information (Figure. 4(a)). To achieve good results, neurons tend to specialize in certain features, resulting in stronger monosemanticity (Figure. 4(b)).
@@
Similarly, we can analogize large models to adults, allowing them to learn and master solving skills related to integration (Figure. 4(c)). This is because adults have accumulated knowledge and have a more mature brain development, similar to the huge training volume and parameter size of large models. In this process, adults gradually reduce the reliance on rote memorization and improve their reasoning abilities for new problems (Figure. 4(d)). This can explain why large models have fewer monosemantic neurons and stronger generalization abilities in complex tasks.
@@ B.1. Artificial Neural Networks and the Increasing Scale
After the Transformer was proposed and validated as effective in various areas (Vaswani et al., 2017), both its scale and performance have seen significant growth over the years (Devlin et al., 2019; Liu et al., 2021). In recent years, architectures based on the Transformer have achieved great success at extremely large scales (Floridi and Chiriatti, 2020; Touvron et al., 2023). Investigating the underlying effective properties created by the increase in scale is a highly demanded research direction.
@@ B.2. Mechanistic Interpretability
B.3. Debate on the Existence of Emergence
@@
Appendix C Proof of Sec. 3
C.1. Proof of Lemma. 3.2
@@
(21) $=$ $\frac{b\left(\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{b}\left(h_{(i+m)}-\mu^{\prime}_{b}\right)^{2}-\sum_{i=1}^{b}\left[2(z^{[m+i]}-\mu^{\prime}_{b})\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$
@@
where the third terms in Equation. (21) and Equation. (22) equal to zero as $\sum_{i=1}^{b}(z^{[m+i]}-\mu^{\prime}_{b})=\sum_{i=1}^{b}(z^{[m+i]})-b\mu^{\prime}_{b}=0,\sum_{i=1}^{m}(z^{[i]}-\mu_{m})=\sum_{i=1}^{m}(z^{[i]})-m\mu_{m}=0$ and other terms are constant factors. ∎
C.2. Proof and Discussion of Lemma. 3.3
@@
Here we focusing on $z-$related parts of Equation. (25) and (26). We denote parameters of $f_{1}$ as $\theta\in\mathbf{W}$ and rewrite Equation. (26) for gradient calculation and parameter update:
(27) $z=$ $g(\theta)+\bar{h}$
@@
where $\frac{\partial t}{\partial\theta}=\nabla g(\theta)$ as the gradient of Equation. (28) and $\frac{\partial(z^{\prime}-z_{a})}{\partial t}=\frac{\partial(-z+z_{a}+\bar{z}-z_{a})}{\partial(z-\bar{z})}=-1$. $\nabla f(\delta)$ denotes $\frac{\partial{\mathcal{L}(f_{2}(\mathbf{z}^{\prime}))}}{\partial(z^{\prime}-z_{a})}.$ According to our assumption that $f(\delta)$ the monotonically increases with $|z^{\prime}-z_{a}|$, $\nabla f(\delta)\geq 0$ when $z^{\prime}\geq z_{a}$ so that $z^{\prime}-z_{a}=|z^{\prime}-z_{a}|$, which means $\bar{z}\geq z_{a}$ and $t\leq 0$. Also, we can derive that $t\geq 0$ when $\nabla f(\delta)\leq 0$.
@@
which is greater than 0 when $d_{2}\geq 0,l>0$ or $d_{2}<0,0<l<\frac{2d_{1}}{-d_{2}}$, where $d_{1}=\nabla f(\delta)||\nabla g(\theta)||^{2}$ and $d_{2}=(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$. As $t=z-\bar{z}<0$ under $\nabla f(\delta)\geq 0$, the activation value is smaller than the mean $\bar{h}$. Note that after the update, $z^{*}>z_{o}$. With a small enough learning rate $l$ (i.e., $<\frac{2d_{1}}{-d_{2}}$), $z^{*}$ will be updated towards $\bar{h}$ and deactivated (i.e., $\bar{h}>z^{*}>z_{o}$)
@@
where Equation. (40) is valid as $g$ is the $z$-related part of model $f$ and has a Lipschitz continuous gradient. Equation. (40) is smaller than 0 when learning rate $0<l<(-\nabla f(\delta))\frac{2}{L}$, which ensures that updated output $z^{*}$ is smaller than original activated value $z_{o}$, and thus close to $\bar{z}$ as $t=z-\bar{z}\geq 0$. Together with the case when $\nabla f(\delta)\geq 0$, we show that $z$ is always pushed to mean $\bar{z}$ and deactivated.
@@
Appendix D Implementation Details
D.1. Implementation Details of MS $\phi$
Algorithm 1 Monosementicity Scale Computing with Needed Variables
  if $c_{t}<n_{f}$ then
  else
  end if
  RETURN: Monosementicity Scale of inputs $\phi({\mathbf{h}})$
Algorithm 2 Monosementicity-based Emergence Learning
  if $c_{t}<l_{s}$ then
     Calculate the MS for each input value
$\phi(h_{i}^{[m+j]})=\frac{(h_{i}^{[m+j]}-{\mu_{i}}_{m})^{2}}{{\upsilon_{i}}_{m}+\epsilon\sum_{k=1}^{n}{\upsilon_{k}}_{m}/n}\text{ for }i\in[1,n],j\in[1,b].$
  else
     $\mathbf{z}^{\prime}=\mathbf{z}$
  end if
  Update sample mean $\{{\mu_{i}}_{m}\}_{i=1}^{n}$ and variance $\{{\upsilon_{i}}_{m}\}_{i=1}^{n}$ using Algorithm. 1.
  RETURN: Adjusted layer of neurons $\mathbf{z}^{\prime}$
D.2. Implementation Details of MEmeL
As described in Algorithm. 1, the statistical variables are calculated online. However, due to the limited number of samples at the beginning of training, the estimation can be unstable. Additionally, if the estimation of $S^{2}$ is extremely small, it may cause overflow during calculation. To improve the robustness of training, we introduce a late start step $l_{s}$ and a variance compensation in the denominator of the metric calculation in Equation. (7). As outlined in Algorithm. 2, when a batch of data arrives, we only update neurons with RD if the current step is greater than $l_{s}$. The calculation of MS is also adjusted by incorporating the mean of variances of other neurons, weighted by a small value $\epsilon$ to prevent overflow.

### v3

@@ 4.1. Experimental Setup
Table 2. Results on ImageNet-1k dataset, where 3 sizes of Swin-Transformer pretrained on ImageNet-22k are used as backbones. The metric used is top-1 accuracy, where a higher value indicates better performance. The best results are indicated in bold font.
Model Swin-T Swin-S Swin-B
@@
Original 80.9 83.2 85.1
Naive (a) 81.0 83.4 84.6
Naive (b) 81.0 83.4 85.1
MEmeL 81.1 83.4 85.1
MEmeL-Tune 81.1 83.5 85.2
Table 3. Results on HKO-7 dataset. We initially trained a ConvGRU model for 20k steps to create the base model. The metrics used are B-MSE and B-MAE, where a smaller value indicates better performance. The best results are in bold fonts. We repeated each experiment three times and reported the average scores.
Model B-MAE B-MSE
Original 1003.41 309.96
Naive (a) 1003.56 309.83
Naive (b) 1003.40 310.13
MEmeL 1003.25 309.94
MEmeL-Tune 998.81 298.16
@@ 4.1.1. Data Sets, Base Models, and Tasks
The configuration of above base models generally follow their original settings (see more details in Appendix E.1). The metrics can be found in Appendix E.2. We emphasize that our method is a general module applicable to models of any scale. After applying our module, the effectiveness of MEmeL can be evaluated by comparing it with the original model.
@@ 4.1.2. Hyper-parameter Setting
At the beginning of the experiment design, we tend to focus on the performance improvement brought by introducing reverse deactivation into milestone models. To fairly validate the influence of the proposed MEmeL with other deactivation approaches (Naive (a) and (b)), we deactivate the neuron with only top-$1$ MS in each batch (recorded as $\rm{MEmeL}$ vs. $\rm{Naive(a)}$ and $\rm{(b)}$). During further discussion, people raised interest about the potential of tuning and we also display the results with tuned hyper-parameters (recorded as $\rm{MEmeL-Tune}$). See more details in Appendix E.3.
@@ 4.1.3. Implementation
Our experiment is conducted on 4 V100 GPU cards. All the codes are implemented in PyTorch (Adam et al., 2017), which are available through the link https://github.com/dominatorX/MEmeL-code.
@@ 4.2. Main Experiment Result
Table 1, Table 2, and Table 3 demonstrate the performance of MEmeL incorporated with BERT, Swin-Transformer, and ConvGRU, respectively. The “Original” method denotes the raw model, “Naive (a)” and “Naive (b)” are corresponding to the original model incorporated with naive inhibition methods (see details in Equation (10)). By comparing the original method and the method enhanced with MEmeL, we can see that the MEmeL (especially MEmeL-Tune) achieves better or comparable results on different tasks, different data sets, and different base models. We also conduct the paired t-test on all three datasets to verify that the improvements are statistically significant (see Appendix F.2). This validates that neural networks can achieve better results by deactivating monosemantic neurons. By comparing Naive (a), (b) with MEmeL (and MEmeL-Tune), we can easily observe that naive deactivation methods may intensify the monosemanticity of neural networks, which may lead to inferior performance improvement as discussed in subsection 3.2.1.
@@ 4.3. Case Study on Deactivating Monosemantic Neurons
To validate that the monosemanticity is indeed inhibited by our reverse deactivation, we conducted experiments on the ImageNet-1k dataset using the Swin-Transformer model as shown in Table 4. We collected 10k samples for 4 different settings, where the modification of $z$ was done using Original (no modification), and two naive methods (a) and (b), and Reverse Deactivation in subsection 3.2.2.
@@ 4.4. Potential and Limitation of MEmeL
According to our hypothesis, MEmeL induces the model to accumulate general and abstract functionality instead of monosemanticity for a specific task, which is consistent with the goal of per-taining. Although MEmeL achieves good results during fine-tuning (demonstrated at Main Experiments in subsection 4.2), the improvement is expected to be even greater when it is applied to the pre-training phase.
The main obstacle is the high computational resource cost required for pre-taining. Currently, we have only completed validation on relatively small precipitation forecasting datasets. In Table 5, the model pre-trained with MEmeL (P-MEmeL) outperforms the one without it (P-Original). Based on these two models, the same finetuning process is conducted without MEmeL. Using MEmeL during pretraining improves B-MAE by 0.18% and B-MSE by 1.29% (T-MEmeL). In contrast, using MEmeL during finetuning improves B-MAE by 0.02% B-MSE by 0.01% (MEmeL in Table 2). The results validate our hypothesis.
@@ 5. Conclusion
Acknowledgements.
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Guangdong Province Science and Technology Plan Project 2023A0505030011, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant and HKUST-Webank joint research lab grants.
@@ Appendix A An Example Analogy that Helps Illustrate Our Motivation
From the perspective of scale differences, when the model is small, the limited number of neurons makes it difficult to abstract and analyze the input, and to provide functional support for integrating a large amount of scattered information (Figure 5(a)). To achieve good results, neurons tend to specialize in certain features, resulting in stronger monosemanticity (Figure 5(b)).
@@
Similarly, we can analogize large models to adults, allowing them to learn and master solving skills related to integration (Figure 5(c)). This is because adults have accumulated knowledge and have a more mature brain development, similar to the huge training volume and parameter size of large models. In this process, adults gradually reduce the reliance on rote memorization and improve their reasoning abilities for new problems (Figure 5(d)). This can explain why large models have fewer monosemantic neurons and stronger generalization abilities in complex tasks.
@@ B.1. Artificial Neural Networks and the Increasing Scale
After the Transformer was proposed and validated as effective in various areas (Vaswani et al., 2017), both its scale and performance have seen significant growth over the years (Devlin et al., 2019; Liu et al., 2021). In recent years, architectures based on the Transformer have achieved great success at extremely large scales (Floridi and Chiriatti, 2020; AI, 2023). Investigating the underlying effective properties created by the increase in scale is a highly demanded research direction.
@@ B.2. Mechanistic Interpretability
B.3. Information Bottleneck
As an important method to explain the deep learning mechanism, the Information bottleneck (IB) provides a compression view for deep learning and emphasizes forgetting together with information retaining (Tishby and Zaslavsky, 2015). For large-scale models, in addition to understanding how much and what information the model retains, understanding how the model manipulates the retained information is also important. Our method encourages the model to organize information more abstractly and distributed into multiple neurons, which is positive to performance supported by existing analysis with biological analogies.
For example, we hope to punish each neuron if it only provides information for a specific feature, rather than considering the total information that a layer of neurons retains and forgets. In other words, we emphasize the way neurons manipulate the retained information (distributed or one-to-one), while IB emphasizes the total amount of information that different layers of neurons record at different training stages (Saxe et al., 2018).
As the scale of neural networks continues to grow, effective information processing and reasoning have become as important a need as information compression. In addition to what information should be squeezed through the bottleneck, how to process the squeezed information is also important and is what our paper concerns (Natalie, 2018). Our method is a promising direction to enrich deep learning understanding together with IB, especially when large model application requires more and more powerful reasoning ability.
B.4. Debate on the Existence of Emergence
@@
Algorithm 1 Monosementicity Scale Computing with Needed Variables
  if $c_{t}<n_{f}$ then
  else
  end if
  RETURN: Monosementicity Scale of inputs $\phi({\mathbf{z}})$
Algorithm 2 Monosementicity-based Emergence Learning
  if $c_{t}<l_{s}$ then
     Calculate the MS for each input value
$\phi(z_{i}^{[m+j]})=\frac{(z_{i}^{[m+j]}-{\mu_{i}}_{m})^{2}}{{\upsilon_{i}}_{m}+\epsilon\sum_{k=1}^{n}{\upsilon_{k}}_{m}/n}\text{ for }i\in[1,n],j\in[1,b].$
  else
     $\mathbf{z}^{\prime}=\mathbf{z}$
  end if
  Update sample mean $\{{\mu_{i}}_{m}\}_{i=1}^{n}$ and variance $\{{\upsilon_{i}}_{m}\}_{i=1}^{n}$ using Algorithm 1.
  RETURN: Adjusted layer of neurons $\mathbf{z}^{\prime}$
Appendix C Implementation Details
C.1. Implementation Details of MS $\phi$
C.2. Implementation Details of MEmeL
As described in Algorithm 1, the statistical variables are calculated online. However, due to the limited number of samples at the beginning of training, the estimation can be unstable. Additionally, if the estimation of $S^{2}$ is extremely small, it may cause overflow during calculation. To improve the robustness of training, we introduce a late start step $l_{s}$ and a variance compensation in the denominator of the metric calculation in Equation (7). As outlined in Algorithm 2, when a batch of data arrives, we only update neurons with RD if the current step is greater than $l_{s}$. The calculation of MS is also adjusted by incorporating the mean of variances of other neurons, weighted by a small value $\epsilon$ to prevent overflow.
Appendix D Proof of Section 3
D.1. Proof of Lemma 3.2
@@ Appendix D Proof of Section 3
(21) $=$ $\frac{b\left(\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right)^{2}+\sum_{i=1}^{b}\left(z_{(i+m)}-\mu^{\prime}_{b}\right)^{2}-\sum_{i=1}^{b}\left[2(z^{[m+i]}-\mu^{\prime}_{b})\frac{m(\mu_{m}-\mu_{b}^{\prime})}{m+b}\right]}{m+b-1}$
@@
where the third terms in Equation (21) and Equation (22) equal to zero as $\sum_{i=1}^{b}(z^{[m+i]}-\mu^{\prime}_{b})=\sum_{i=1}^{b}(z^{[m+i]})-b\mu^{\prime}_{b}=0,\sum_{i=1}^{m}(z^{[i]}-\mu_{m})=\sum_{i=1}^{m}(z^{[i]})-m\mu_{m}=0$ and other terms are constant factors. ∎
D.2. Proof and Discussion of Lemma 3.3
@@
Here we focusing on $z-$related parts of Equation (25) and (26). We denote parameters of $f_{1}$ as $\theta\in\mathbf{W}$ and rewrite Equation (26) for gradient calculation and parameter update:
(27) $z=$ $g(\theta)+\bar{z}$
@@
where $\frac{\partial t}{\partial\theta}=\nabla g(\theta)$ as the gradient of Equation (28) and $\frac{\partial(z^{\prime}-z_{a})}{\partial t}=\frac{\partial(-z+z_{a}+\bar{z}-z_{a})}{\partial(z-\bar{z})}=-1$. $\nabla f(\delta)$ denotes $\frac{\partial{\mathcal{L}(f_{2}(\mathbf{z}^{\prime}))}}{\partial(z^{\prime}-z_{a})}.$ According to our assumption that $f(\delta)$ the monotonically increases with $|z^{\prime}-z_{a}|$, $\nabla f(\delta)\geq 0$ when $z^{\prime}\geq z_{a}$ so that $z^{\prime}-z_{a}=|z^{\prime}-z_{a}|$, which means $\bar{z}\geq z_{a}$ and $t\leq 0$. Also, we can derive that $t\geq 0$ when $\nabla f(\delta)\leq 0$.
@@
which is greater than 0 when $d_{2}\geq 0,l>0$ or $d_{2}<0,0<l<\frac{2d_{1}}{-d_{2}}$, where $d_{1}=\nabla f(\delta)||\nabla g(\theta)||^{2}$ and $d_{2}=(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$. As $t=z-\bar{z}<0$ under $\nabla f(\delta)\geq 0$, the activation value is smaller than the mean $\bar{z}$. Note that after the update, $z^{*}>z_{o}$. With a small enough learning rate $l$ (i.e., $<\frac{2d_{1}}{-d_{2}}$), $z^{*}$ will be updated towards $\bar{z}$ and deactivated (i.e., $\bar{z}>z^{*}>z_{o}$)
@@
where Equation (40) is valid as $g$ is the $z$-related part of model $f$ and has a Lipschitz continuous gradient. Equation (40) is smaller than 0 when learning rate $0<l<(-\nabla f(\delta))\frac{2}{L}$, which ensures that updated output $z^{*}$ is smaller than original activated value $z_{o}$, and thus close to $\bar{z}$ as $t=z-\bar{z}\geq 0$. Together with the case when $\nabla f(\delta)\geq 0$, we show that $z$ is always pushed to mean $\bar{z}$ and deactivated.
@@
Table 6. Validation results on GLUE Test datasets. The settings are the same with Table 1. All metrics are the larger the better with best results in bold font.
Model MNLI-(M/MM) QQP QNLI SST-2 CoLA STS-B MRPC RTE Average
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 68.2 80.2
MEmeL-T 84.9/83.7 71.5 90.5 93.9 52.1 85.9 88.7 68.1 79.9
Table 7. Validation results on ImageNet-1k dataset. See Table 2 for detailed settings. The metric is the higher the better. The best results are indicated in bold font.
Model Swin-T Swin-S Swin-B
Size 28M 50M 88M
MEmeL 81.1 83.4 85.1
MEmeL-T 81.1 83.5 85.1
Table 8. Results on HKO-7 dataset. The settings are the same with Table 3. The metrics are the smaller the better with the best results in bold fonts.
Model B-MAE B-MSE
MEmeL 1003.25 309.94
MEmeL-T 1002.46 309.52